# How Generalizable Is My Behavior Cloning Policy? A Statistical Approach to Trustworthy Performance Evaluation

### v1

@@ Abstract
With the rise of stochastic generative models in robot policy learning, end-to-end visuomotor policies are increasingly successful at solving complex tasks by learning from human demonstrations. Nevertheless, since real-world evaluation costs afford users only a small number of policy rollouts, it remains a challenge to accurately gauge the performance of such policies. This is exacerbated by distribution shifts causing unpredictable changes in performance during deployment. To rigorously evaluate behavior cloning policies, we present a framework that provides a tight lower-bound on robot performance in an arbitrary environment, using a minimal number of experimental policy rollouts. Notably, by applying the standard stochastic ordering to robot performance distributions, we provide a worst-case bound on the entire distribution of performance (via bounds on the cumulative distribution function) for a given task. We build upon established statistical results to ensure that the bounds hold with a user-specified confidence level and tightness, and are constructed from as few policy rollouts as possible. In experiments we evaluate policies for visuomotor manipulation in both simulation and hardware. Specifically, we (i) empirically validate the guarantees of the bounds in simulated manipulation settings, (ii) find the degree to which a learned policy deployed on hardware generalizes to new real-world environments, and (iii) rigorously compare two policies tested in out-of-distribution settings. Our experimental data, code, and implementation of confidence bounds are open-source.  [^1] https://github.com/TRI-ML/stochastic_verification
    https://github.com/TRI-ML/binomial_CIs
@@ I Introduction
In this paper we focus on evaluating robot policies obtained through the behavior cloning (BC) framework of robot learning. BC policies, such as diffusion policies [1], have recently advanced the state-of-the-art (SOTA) in visuomotor policy synthesis, particularly in manipulation tasks [1, 2, 3, 4, 5]. BC is attractive because it avoids the significant challenges of the sim-to-real gap, which impede the transfer of reinforcement learning policies and other policy synthesis techniques to real-world robots [6]. In BC, humans give demonstrations, often directly on the robot hardware through teleoperation. Then, a policy is learned based on these demonstrations. This procedure removes the need for a simulation model, avoiding the sim-to-real gap.
@@
However, without an accurate simulation model, evaluation of BC policies relies on real-world tests. In robotics research it is common to evaluate these policies using fewer than $50$ policy rollouts, recording the empirical success rate or average reward (see, e.g., [1, 2, 7]). However, with such small sample sizes, it can be difficult to interpret the significance of the recorded results. In addition, measuring average performance can be insufficient for applications with safety and reliability requirements.
@@
Our proposed method is to compute worst-case bounds on the performance of policy based on the results of a small number of policy rollouts. Worst-case bounds are useful for determining a policy exceeds some performance threshold rather than determining it does not. Our approach can specify the fewest number of policy rollouts required to obtain a user-specified confidence and tightness for a bound on the performance metric. Confidence is the probability (fraction of random outcomes) for which the bound holds. [^2] Sometimes also referred to as the coverage of the bound. Tightness quantifies how close the bound tends to be to the true (but unknown) quantity of interest. A bound which is tighter provides a better estimate of the unknown quantity while a bound which is higher confidence provides a better chance that the estimate is conservative (rather than optimistic). In Fig. 2 we give a visual interpretation of how confidence and tightness relate to the distribution of a lower confidence bound for an unknown probability of success. Achieving higher confidence and tighter bounds on policy performance comes at the cost of more policy rollouts.
@@
The bounds we use require no knowledge of the underlying probability distributions in the policy, training data, robot hardware, or environment physics. The confidence and tightness of these bounds do not depend on the dimensionality of the robot state or observation. We only require that the stochasticity of the policy, robot, and environment is stationary between rollouts so that the measured performance of the policy rollouts can be treated as independent and identically distributed (i.i.d.) random variables. This assumption is frequently implicit during BC training, and, depending on the task, can be effectively managed during evaluation.
@@
In this paper we demonstrate how to construct worst-case bounds on performance when performance is measured by (i) a binary success/failure metric, and (ii) a continuous metric which measures the reward accumulated along a trajectory. For the binary metric, we seek a lower bound on the unknown probability of task success. We revisit a rarely-used bound from the statistics literature, which is a tighter version of the more common Clopper-Pearson bound [8]. Our bound exactly achieves a user-specified confidence level and is as tight as possible. This is enabled by a novel computational method based on mixed monotonic optimization [9]. For the continuous metric, the true distribution of reward is entirely unknown. We propose to use bounds for the cumulative distribution function (CDF) of reward. The CDF fully characterizes the reward distribution, allowing flexibility for the user to analyze performance, e.g., in the tails, at any quantile, or in the mean. We use a tighter version of the Dvoretzky–Kiefer–Wolfowitz (DKW) bound [10, 11] for this purpose. In both the binary and continuous cases, the user specifies the confidence level and the tightness of the bound. Since real-world evaluation is costly, the bounds we employ meet these specifications exactly using the minimum number of policy rollouts. This is in contrast to the more common Clopper-Pearson and DKW bounds.
@@
Finally, in our experiments we are primarily interested in applying our framework to investigate the generalization of BC policies to out-of-distribution (OOD) settings. Because human intuition is often unreliable for predicting which OOD settings will be harmful to the policy, we believe this is an application where statistical evaluation methods shine, giving roboticists well-defined confidence levels for the results of their experiments. We demonstrate our bounds in simulation and real-world robotic manipulation settings. We empirically compare our bounds to more common bounds from the literature, showing improved tightness, particularly for few policy rollouts. We also show how our bounds can be used to compare the performance of two policies, and to help decide which policy is superior. Finally, in hardware, we bound the performance of a SOTA visuomotor diffusion policy [1] in generalizing to new environments. An early version of these results appeared in a workshop [12]. Here we substantially expand the analysis and simulation results, add hardware experiments, and add an example use-case for comparing policies. We summarize our contributions as follows:
@@ II-A Formal Evaluation of Learned Policies
Formal methods applied to robots with learned components can be used to evaluate open-loop policy requirements (such as state-dependent action constraints [13]) and closed-loop system requirements (such as reach-avoid constraints [14, 15, 16, 17]). These methods typically have strict limitations on the structure and complexity of the robot policy and environment model. In addition, these methods are most commonly employed in settings with deterministic environments or with bounded (set-based) uncertainties. In settings where these requirements are met, formal methods can evaluate rich closed-loop behaviors by finding certificates such as invariant sets [15, 16], Lyapunov functions [18, 19], and contraction metrics [20, 21]. However, in this paper we do not assume access to an environment model or any particular policy structure, precluding the use of formal evaluation methods. When systems become too complex for formal evaluation, black-box validation methods are often used [22]. Statistical testing, discussed next, is one flavor of black-box validation that provides probabilistic guarantees.
@@ II-B Statistical Evaluation of Learned Policies
When constructing statistical bounds, there are trade-offs between confidence, tightness, and sample size. Related work tends to focus on the trade-off between confidence and number of samples, neglecting the important role that tightness plays in interpreting the bound. Placing statistical bounds on the performance of a learning-enabled robot from few policy rollouts is explored in [23, 24, 25, 26]. However, these works only bound particular scalar quantities such as the expected value, value-at-risk, and conditional value-at-risk of the performance distribution. We place bounds on the CDF that consider all aspects of the performance distribution. For example, policies with appealing expected performance may have unacceptable long tails of poor performance. Further, these papers do not quantify bound tightness and their bounds can require more policy rollouts than necessary to achieve the specified confidence.
Conformal prediction has been used for rigorous uncertainty quantification of robot perception components [27, 28, 29, 30] and policies [31, 32]. This research has enabled policies to attain a specified probability of task success, such as avoiding collisions or making language-based plans. However, like the other approaches in this section, (i) by bounding a quantile of the score distribution, conformal prediction bounds also ignore tails of distributions, (ii) tightness of bounds from conformal prediction is often not quantified, and (iii) conformal prediction bounds may require more samples than necessary to achieve the desired confidence level [33, 34]. In this paper we bound the entire distribution of performance, quantify bound tightness, and ensure that the number of samples used is minimal.
Confidence bounds for unknown success probabilities and CDFs are well studied with the Wald and Clopper-Pearson intervals being the most common for success probabilities [35, 8] and with the DKW inequality being the most common for CDFs [10, 11]. However, these bounds are not optimally tight and thus may require more policy rollouts than necessary. In Section IV and Section V, we discuss less common yet well-established confidence bounds which optimally trade-off between confidence, tightness, and number of samples. Lastly, in Section VI-C we compare the success rates of two policies using the bounds in this paper. Common approaches which test for differences in success rates include Fisher’s and Barnard’s tests [36]. However, while such tests determine which policy performs better, they do not return confidence intervals for the individual success rates (as we do), but rather provide confidence intervals for the odds ratio. Our view is that in practice it is not enough to know which policy performs better, but it is also important to know the performance of each policy individually.
@@ III Distributional Bounds
$\overline{F}(x)$ is constructed based on the observed performance of some number of policy rollouts. Since the outcome of policy rollouts is random, $\overline{F}(x)$ is also random, leading to the bound holding probabilistically as in Eq. 1. In light of the above discussion on partial ordering of CDFs, $\overline{F}(x)$ is a worst-case bound on the distribution of performance with confidence $1-\alpha$. Next, we give worst-case bounds for the both binary and continuous performance metrics. In the case of a binary performance metric, an upper bound on the CDF is simply a lower bound on the probability of success.
@@ IV Confidence Bounds - Binary Metric
To find lower bounds on a policy’s success rate in a new environment, we treat the result of each policy rollout as a Bernoulli random variable $X$, where $X=1$ denotes a success and $X=0$ denotes a failure. Then given the results of $n$ i.i.d. policy rollouts, we construct a lower confidence bound $\underline{p}$. The number of observed successes follows a binomial distribution. Before describing the bound, we first introduce general optimality criteria for lower confidence bounds of an unknown parameter $\theta$.
@@ IV-A Optimality Criteria
Definition 1 (Uniformly Most Accurate, Eq. 3.22 of [37])
@@
and amongst all possible lower bounds $\underline{\theta}^{\prime}(T)$ minimizes
@@
Now, since we know that a UMA lower confidence bound minimizes the integrand of Eq. 5, we can conclude that a UMA lower confidence bound also minimizes expected shortage for all values of $\theta$. Although expected shortage is a useful quantity, it depends on the value of $\theta$, which is unknown. To avoid any assumptions on $\theta$, we measure tightness according to maximum expected shortage (MES),
$\textup{MES}=\max_{\theta}\ \mathbb{E}_{\theta}[\textup{shortage}].$ (6)
@@ IV-B Optimal Binomial Bounds
The bisection method can be used to solve $F_{p^{*}}(t)=1-\alpha$, as $F_{p}(t)$ is monotonically decreasing in $p$. Taking limits $p\rightarrow 0^{+}$ and $p\rightarrow 1^{-}$, one can show that $F_{p}(t)=1-\alpha$ has no solution when $t<1-\alpha$ and $t>n+1-\alpha$, we thus set $\underline{p}$ to either zero or one in these cases. This bound is well-established (see Example 3.5.2 of [37]). The Clopper-Pearson bound can be obtained by setting $U=0$ in the test statistic.
@@
In this form, expected shortage is increasing in $p_{1}$ and decreasing in $p_{2}$; this property is known as mixed monotonicity. Note that for $p_{1}=p_{2}$ we recover the original function. Mixed monotonic functions can be globally optimized using the branch and bound procedure given in [9]. This approach is the first tractable method that certifiably solves the optimization needed to compute MES. We provide an open-source implementation of this UMA bound at     https://github.com/TRI-ML/binomial_CIs.
Now, given two of the following: (i) confidence level ($1-\alpha$), (ii) tightness (MES), (iii) number of policy rollouts ($n$), we can determine the third. In Fig. 3 we show how these quantities vary with one another for our bound and the Clopper-Pearson bound. Since the bound we use is UMA, it always has lower MES than Clopper-Pearson. MES computation for Clopper-Pearson is also novel. We derive the appropriate form of Eq. 12 and compute the corresponding MES using mixed monotonic programming. Lastly, one can compute a lower confidence bound on the failure probability using the same procedure, but constructing the test statistic $T$ with failure counts rather than successes.

### v2

@@ Abstract
With the rise of stochastic generative models in robot policy learning, end-to-end visuomotor policies are increasingly successful at solving complex tasks by learning from human demonstrations. Nevertheless, since real-world evaluation costs afford users only a small number of policy rollouts, it remains a challenge to accurately gauge the performance of such policies. This is exacerbated by distribution shifts causing unpredictable changes in performance during deployment. To rigorously evaluate behavior cloning policies, we present a framework that provides a tight lower-bound on robot performance in an arbitrary environment, using a minimal number of experimental policy rollouts. Notably, by applying the standard stochastic ordering to robot performance distributions, we provide a worst-case bound on the entire distribution of performance (via bounds on the cumulative distribution function) for a given task. We build upon established statistical results to ensure that the bounds hold with a user-specified confidence level and tightness, and are constructed from as few policy rollouts as possible. In experiments we evaluate policies for visuomotor manipulation in both simulation and hardware. Specifically, we (i) empirically validate the guarantees of the bounds in simulated manipulation settings, (ii) find the degree to which a learned policy deployed on hardware generalizes to new real-world environments, and (iii) rigorously compare two policies tested in out-of-distribution settings. Our experimental data, code, and implementation of confidence bounds are open-source.  [^1] Project Page: https://tri-ml.github.io/stochastic_verification
Paper Repository: https://github.com/TRI-ML/stochastic_verification
Binomial Confidence Intervals: https://github.com/TRI-ML/binomial_cis
@@ I Introduction
In this paper we focus on evaluating robot policies obtained through the behavior cloning (BC) framework of robot learning. BC policies, such as diffusion policies [1], have recently advanced the state-of-the-art (SOTA) in visuomotor policy synthesis, particularly in manipulation [1, 2, 3, 4, 5]. BC is attractive because it avoids the challenges of the sim-to-real gap, which impede the transfer of policies from reinforcement learning and other techniques to real-world robots [6]. In BC, humans give demonstrations, often directly on the robot hardware through teleoperation. Then, a policy is learned based on these demonstrations. This procedure removes the need for a simulation model, avoiding the sim-to-real gap.
@@
However, without an accurate simulation model, evaluation of BC policies relies on real-world tests. In robotics research it is common to evaluate these policies using fewer than $50$ policy rollouts, recording the empirical success rate or average reward (e.g., [1, 2, 7]). With such small sample sizes, it can be difficult to interpret the significance of the recorded results. In addition, measuring average performance can be insufficient for applications with safety and reliability requirements.
@@
Our proposed method is to compute worst-case bounds on the performance of a policy based on the results of a small number of policy rollouts. In Section III, we define an ordering of distributions of policy performance. A worst-case bound on the performance of a policy is the least-preferable distribution according to this ordering that is consistent with the observed data (up to some confidence level). Worst-case bounds are useful for determining whether a policy exceeds some performance threshold rather than determining it does not. Our approach can specify the fewest number of policy rollouts required to obtain a user-specified confidence and tightness for a bound on the performance metric. Confidence is the probability (fraction of random outcomes) for which the bound holds. [^2] Sometimes also referred to as the coverage of the bound. Tightness quantifies how close the bound tends to be to the true (but unknown) quantity of interest. A bound which is tighter provides a better estimate of the unknown quantity while a bound which is higher confidence provides a better chance that the estimate is conservative (rather than optimistic). In Fig. 2 we give a visual interpretation of how confidence and tightness relate to the distribution of a lower confidence bound for an unknown probability of success. Achieving higher confidence and tighter bounds on policy performance comes at the cost of more policy rollouts.
@@
The bounds we use require no knowledge of the underlying probability distributions in the policy, training data, robot hardware, or environment physics. The confidence and tightness of these bounds do not depend on the dimensionality of the robot state or observation. We only require performance measurements to be i.i.d. which we further discuss in Section III-B.
@@
In this paper we demonstrate how to construct worst-case bounds on performance when performance is measured by (i) a binary success/failure metric, and (ii) a continuous metric which measures the reward accumulated along a trajectory. For a binary metric, we seek a lower bound on the unknown probability of task success. We revisit a rarely-used bound from the statistics literature, a tighter version of the more common Clopper-Pearson bound [8]. Our bound exactly achieves a user-specified confidence level and is as tight as possible. This is enabled by a novel computational method based on mixed monotonic optimization [9]. For a continuous metric, the true distribution of reward is entirely unknown. We propose to use bounds for the cumulative distribution function (CDF) of reward. The CDF fully characterizes the reward distribution, allowing flexibility for the user to analyze performance, e.g., in the tails, at any quantile, or in the mean. We use a tighter version of the Dvoretzky–Kiefer–Wolfowitz (DKW) bound [10, 11] for this purpose. In both the binary and continuous cases, the user specifies the confidence level and the tightness of the bound. Since real-world evaluation is costly, the bounds we employ meet these specifications using the minimum number of policy rollouts. This is in contrast to the more common Clopper-Pearson and DKW bounds.
@@
Finally, in our experiments we are primarily interested in applying our framework to investigate the generalization of BC policies to out-of-distribution (OOD) settings. We accomplish this by directly testing in the OOD settings. Because human intuition is often unreliable for predicting which OOD settings will be harmful to the policy, we believe this is an application where statistical evaluation methods shine, giving roboticists well-defined confidence levels for the results of their experiments. We demonstrate our bounds in simulation and real-world manipulation settings. We compare our bounds to more common bounds from the literature, showing improved tightness, particularly for few policy rollouts. We also show how our bounds can be used to compare the performance of two policies. Finally, in hardware, we bound the performance of a SOTA visuomotor diffusion policy [1] in new environments. An early version of these results appeared in a workshop [12]. Here we substantially expand the analysis and simulation results, add hardware experiments, and add a policy comparison example. Our contributions are as follows:
@@ II-A Formal Evaluation of Learned Policies
Formal methods applied to robots with learned components can be used to evaluate open-loop policy requirements (such as state-dependent action constraints [13]) and closed-loop system requirements (such as reach-avoid constraints [14, 15, 16, 17]). These methods typically have strict limitations on the structure and complexity of the policy and environment model. In addition, these methods are most commonly employed in settings with deterministic environments or with bounded (set-based) uncertainties. In settings where these requirements are met, formal methods can evaluate rich closed-loop behaviors by finding certificates such as invariant sets [15, 16], Lyapunov functions [18, 19], and contraction metrics [20, 21]. However, in this paper we do not assume access to an environment model or any particular policy structure, precluding the use of formal evaluation methods. When systems become too complex for formal evaluation, black-box validation methods are used [22]. Statistical testing, discussed next, is one flavor of black-box validation that provides probabilistic guarantees.
@@ II-B Statistical Evaluation of Learned Policies
When constructing statistical bounds, there are trade-offs between confidence, tightness, and sample size. Related work tends to focus on the trade-off between confidence and number of samples, neglecting the important role that tightness plays in interpreting the bound. Placing statistical bounds on the performance of learned policies from few rollouts is explored in [23, 24, 25, 26]. However, these works only bound particular scalar quantities such as the expected value, value-at-risk, and conditional value-at-risk of the performance distribution. We place bounds on the CDF that consider all aspects of the performance distribution. For example, policies with appealing expected performance may have unacceptable long tails of poor performance. Further, these papers do not quantify bound tightness and their bounds can require more policy rollouts than necessary to achieve the specified confidence.
Conformal prediction has been used for uncertainty quantification of robot perception components [27, 28, 29, 30] and policies [31, 32]. This research has enabled policies to attain a specified probability of task success, such as avoiding collisions or making language-based plans. However, like the other approaches in this section, (i) by bounding a quantile of the score distribution, conformal prediction bounds also ignore tails of distributions, (ii) tightness of bounds from conformal prediction is often not quantified, and (iii) conformal prediction may require more samples than necessary to achieve a desired confidence [33, 34]. In this paper we bound the entire distribution of performance, quantify bound tightness, and ensure that the number of samples used is minimal.
Confidence bounds for unknown success probabilities and CDFs are well studied with the Clopper-Pearson bound being common for success probabilities [8] and with the DKW inequality being common for CDFs [10, 11]. However, these bounds are loose and may require more policy rollouts than necessary. In Section IV and Section V, we discuss less common yet well-established confidence bounds which optimally trade-off between confidence, tightness, and number of samples. Lastly, in Section VI-C we compare the success rates of two policies using the bounds in this paper. Common approaches which test for differences in success rates include Fisher’s and Barnard’s tests [35]. These are the standard methods for performing A/B tests to distinguish success rates. However, while such tests determine which policy performs better, they do not return confidence intervals for the individual success rates (as we do), but rather provide confidence intervals for the odds ratio. The drawback is that the odds ratio lacks information about absolute performance, i.e., pairs of policies with low success rates and with high success rates can have the same odds ratio. Our view is that it is not enough to know which policy performs better, but it is also important to know the performance of each policy individually.
@@ III-A Bounds
$\overline{F}(x)$ is constructed based on the observed performance of some number of policy rollouts. Since the outcome of policy rollouts is random, $\overline{F}(x)$ is also random, leading to the bound holding probabilistically as in Eq. 1. In light of the above discussion on partial ordering of CDFs, $\overline{F}(x)$ is a worst-case bound on the distribution of performance with confidence $1-\alpha$. In the case of a binary performance metric, an upper bound on the CDF is simply a lower bound on the success rate.
III-B Assumptions
The methods we use in Sections IV and V to construct $\overline{F}(x)$ rely on i.i.d. performance measurements.
Assumption 1.
The performance measurements from the policy rollouts are i.i.d. random variables.
That the performance measurements are random variables is ensured by a stochastic policy and/or stochastic environment. To be independent, the outcome of one policy rollout must have no influence on the outcome of another. To be identically distributed, there must be no distribution shift between rollouts. For binary metrics, 1 ensures the performance will follow a Bernoulli distribution with unknown probability of success. For continuous metrics, the assumption allows for the performance to follow any distribution (including discrete and mixed continuous-discrete distributions).
To avoid violating 1, a testing plan should be developed before collecting any policy rollouts, documenting the sampling rule for initial conditions and the number of policy rollouts. Practices which violate 1 include
  • running policy rollouts until enough favorable outcomes are observed (in which case confidence sequences would be more appropriate [37]),
  • running policy rollouts in a time-varying environment (e.g., lighting changes between rollouts due to sunset).
For further discussion on misuse of statistical evaluation see [38]. Next, we give worst-case bounds for both the binary and continuous performance metrics.
@@ IV Confidence Bounds - Binary Metric
To find lower bounds on a policy’s success rate in a new environment, we treat the result of each policy rollout as a Bernoulli random variable $X$, where $X=1$ denotes a success and $X=0$ denotes a failure. Then, given the $n$ i.i.d. success/failure measurements, we construct a lower confidence bound $\underline{p}$. The number of observed successes follows a binomial distribution. Before describing the bound, we first introduce optimality criteria for lower confidence bounds.
@@ IV-A Optimality Criteria
Definition 1 (Uniformly Most Accurate, Eq. 3.22 of [36]).
@@
and, amongst all possible lower bounds $\underline{\theta}^{\prime}(T)$, minimizes
@@
Now, since a UMA lower confidence bound minimizes the integrand of Eq. 5, a UMA lower confidence bound also minimizes expected shortage for all values of $\theta$. Although expected shortage is a useful quantity, it depends on the unknown value of $\theta$. To avoid assumptions on $\theta$, we measure tightness according to maximum expected shortage (MES),
[…]
@@ IV-B Optimal Binomial Bounds
Note that since the test statistic is defined as a sum of uniform and binomial random variables, given the same number of observed successes, $t$ will be almost surely distinct, resulting in distinct values for $\underline{p}(t)$. Thus, $\underline{p}$ is known as a randomized confidence bound, and this property is necessary to construct a UMA confidence bound for an unknown success rate.
A bisection search can be used to solve Eq. 9b, as $F_{p}(t)$ is decreasing in $p$. Taking limits $p\rightarrow 0^{+}$ and $p\rightarrow 1^{-}$, one finds $F_{p}(t)=1-\alpha$ has no solution when $t<1-\alpha$ and $t>n+1-\alpha$. This bound is well-established (see Example 3.5.2 of [36]). The Clopper-Pearson bound is obtained by setting $U=0$ in the test statistic.
@@
In this form, expected shortage is increasing in $p_{1}$ and decreasing in $p_{2}$; this property is known as mixed monotonicity. Mixed monotonic functions can be globally optimized [9]. This approach is the first tractable method that certifiably solves the optimization needed to compute MES and is included in our open-source implementation.333    https://github.com/TRI-ML/binomial_cis
Now, given two of the following: (i) confidence level ($1-\alpha$), (ii) tightness (MES), (iii) number of policy rollouts ($n$), we can determine the third. For example, since MES is decreasing in $n$, to determine $n$ given a desired $\alpha$ and MES, one fixes $\alpha$ and computes the smallest $n$ (via a bisection search) which yields an MES below the desired value. In Fig. 3 we show how these quantities vary with one another for our bound and the Clopper-Pearson bound. Since the bound we use is UMA, it always has lower MES. Also novel is the MES computation for Clopper-Pearson. We derive the appropriate form of Eq. 12 and compute MES using mixed monotonic programming. [^4] Further details on this derivation for the Clopper-Pearson bound are in the documentation of the linked binomial_cis Github repository. Lastly, one can compute a lower confidence bound on the failure probability using the same procedure, but with failure counts in the test statistic $T$ rather than success counts.